OpenStudy (anonymous):

a spherical fish bowl has a radius of 21cm. how many ml of water is needed to fill half the bowl?

OpenStudy (gorv):

volume of sphere given by formula \[\frac{ 4 }{ 3 } *\pi* r^{3}\]

OpenStudy (gorv):

to fill half just half the value you get from formula

OpenStudy (anonymous):

a spherical fish bowl has a radius of 21cm. how many ml of water is needed to fill half the bowl?

OpenStudy (anonymous):

for full and for half devide by 2

OpenStudy (gorv):

=(1/2 )*(4/3)*pi* (0.21 m )^3

OpenStudy (gorv):

after calculation convert volume in ml

OpenStudy (alekos):

not quite to fill the bowl you need 2/3 πr^3 cc's

OpenStudy (alekos):

thats 2/3pi*r^3

OpenStudy (anonymous):

for full and for half devide by 2 right

OpenStudy (alekos):

no not quite. If the height of the water level is half the radius of the bowl then the volume is 5/24*pi*r^3

OpenStudy (gorv):

alekos u feel sleepy

OpenStudy (alekos):

V = (pi*h^2)(3r-h)/3

OpenStudy (gorv):

that is a sphere ...

OpenStudy (gorv):

half fill means u have to fill half oh the sphere that is half volume of the sphere to be filled

OpenStudy (alekos):

No a fish bowl is half a sphere

OpenStudy (gorv):

read the Q it is spherical nothing is said about its base

OpenStudy (gorv):

so we will consider it like a ball

OpenStudy (alekos):

It says " a spherical fish bowl" not a "sphere"

OpenStudy (anonymous):

ok i git it

OpenStudy (alekos):

mirzah. have they given you an answer?

OpenStudy (gorv):

okkk how u will find the height

OpenStudy (alekos):

i'm assuming that the height is half the radius because it says "half fill the bowl"

OpenStudy (gorv):

hahahahaha...that means it is full sphere lolllllllll

OpenStudy (alekos):

no it aint

OpenStudy (gorv):

than how u can consider h=r/2

OpenStudy (gorv):

for ur kind info bowl will be flat at base ...so not complete sphere and mouth will also open

OpenStudy (alekos):

because it says "half fill the bowl" which I'm assuming to be a hemi-spherical bowl

OpenStudy (alekos):

doesn't say bowl with flat base

OpenStudy (gorv):

THAN WHAT I SAID IN START

OpenStudy (gorv):

THAT WE HAVE TO CONSIDER IT AS SPHEREICAL

OpenStudy (alekos):

all depends on how you interpret the question i say tomaato and you say tomayto